# Scalable multi-qubit intrinsic gates in quantum dot arrays

### v1

@@ IV.3 Dynamical calibration: general theory
We now present a general theory for the pulsed evolution in the presence of bond inhomogeneity. For a general array of quantum dots with given connection, the intrinsic gate can be decomposed into the product of MQCP gates. The corresponding time evolution reduces to the time evolution in a high dimensional $k$-spaces according the equation
@@
where $\bm{\Delta}=(\Delta_{1},\Delta_{2},\cdots)$ is the vector of effective strength for all the connecting bonds in the array, and $\bm{\theta}=(\theta_{1},\theta_{2},\cdots)$ is the vector of target phase factors. The $\pi$-modulus makes the target gate represents by a high-dimensional lattice of target points. However, the bond strength may differ from ideal values, resulting in unwanted errors in gate implementations. For example, to implement simultaneous multi-qubit controlled-Z gates, all components of $\bm{\theta}$ are equal to $\pi$ and it is desirable that the effective strength $\Delta_{w}$ for all the participating bonds are equal as well. But the actual bond strengths can vary across the full array.
We can overcome the bond inhomogeneity with recurrent single-qubit pulses. This is done by keeping the dots coupled while inserting single-qubit pulses during the evolution. In general, we can write the qubit-frame time evolution operator for the pulsed evolution as
@@
where the pulse $P_{n}$ is applied at time instance $\tau_{0}+\tau_{1}+\cdots+\tau_{n}$ to some vertex point, within the total duration $\tau=\tau_{0}+\tau_{1}+\cdots+\tau_{N}$. The reverse rotation $\mathrm{e}^{\mkern 1.0mu\mathrm{i}\mkern 1.0mu\tau H_{0}}$ in Eq. (69) transforms the lab-frame map to the qubit-frame. The applied pulses are chosen from the Pauli group and they compose into $P_{N}P_{N-1}\cdots P_{1}=Q$. Further introducing an auxiliary set of pulses according to $Q_{n}=P_{n}P_{n-1}\cdots P_{1}$, with $Q_{0}=I$ and $Q_{N}=Q$, we can rewrite the time evolution as
@@
where we have invoked the perturbative assumption $\lVert H_{0}\rVert\gg\lVert H_{\mathrm{ex}}\rVert$ for the approximation in Eq. (70). Since $H_{0}$ involves only Pauli-Z operators $\{\sigma_{j}^{\mathrm{Z}}\}$, conjugation of $H_{0}$ by $Q_{n}$ only selectively flips the signs for the $\sigma_{j}^{\mathrm{Z}}$ operators and the resulting matrix is still diagonal in the computational basis. Hence we can commute and combine terms in Eq. (70) and further derive
@@
The time evolution now comprises of three parts. First, an extra Pauli gate $Q$ is introduced by the applied pulses. It can be corrected by applying another gate later in the circuit, or by simply choosing a pulse sequence that composes into the identity. The second part in Eq. (71) is a set of local phase shifts induced by the $X_{j}$ and $Y_{j}$ conjugations with $\sigma_{j}^{\mathrm{Z}}$. Assuming $Q_{n}=\bigotimes_{j}Q_{nj}$ with $Q_{kj}\in\{I,X,Y,Z\}$ for the $j$th qubit in the array. For a given pulse scheme, the additional phases is explicitly determined by
@@
where $\operatorname{sig}(I)=\operatorname{sig}(Z)=1$ and $\operatorname{sig}(X)=\operatorname{sig}(Y)=-1$. The extra local phase shifts can be easily corrected with virtual-$Z$ gates and we may ignore them for now. The third part in Eq. (71) describes the entangling dynamics of the array, which is a simple product of $N+1$ time-independent stages. Each stage is generated by a Pauli-conjugated grid vector $\Lambda^{(n)}\equiv Q^{\dagger}_{n}\Lambda Q_{n}$ that can be decomposed as the Kronecker sum
@@
By expressing each bond vector as the diagonal matrix $\Lambda_{w}\mathrel{\widehat{=}}\operatorname{diag}(S_{w},T_{w},T_{w},S_{w})$, one can directly verify that $X$ or $Y$ conjugations on either dot exchange the $T_{w}$ and $S_{w}$ tunneling coefficients, while $\Lambda_{w}$ is invariant under $Z$ conjugations. Applying $X$ or $Y$ conjugation to both $j$ and $k$ exchanges the tunneling coefficients twice and leaves the bond vector in its original state. For example, $X_{j}\Lambda_{w}X_{j}=Z_{j}Y_{k}\Lambda_{w}Z_{j}Y_{k}\mathrel{\widehat{=}}\operatorname{diag}(T_{w},S_{w},S_{w},T_{w})$ and $X_{j}Y_{k}\Lambda_{w}X_{j}Y_{k}=\Lambda_{w}$. The exchange of tunneling coefficients effectively results in a reversal of velocity component for bond $w$,
$\Delta_{w}\to-\Delta_{w}.$ (74)
For the entire array of dots, an $X$ or $Y$ gate applied to one vertices point will simultaneously reverse the velocities for all bonds connected to that dot. Combined with the geometry of the array, each $Q_{n}$ specifies the signs for all the bond velocity components. The $\Lambda^{(n)}$ at different stages of evolution are just different combinations of bond velocity components to be integrated with evolution time. This provide a mean to dynamically calibrating the bond strength by properly choosing the time and locations of applied pulses.
@@
To demonstrate the principles of dynamical calibration, we consider the rectangular grid shown in Fig. 7(a). The dots are connected with four inhomogeneous bonds labeled by ‘N’, ‘E’, ‘W’ and ‘S’. Without loss of generality, we assume $\Delta_{\mathrm{N}}<\Delta_{\mathrm{E}}<\Delta_{\mathrm{W}}<\Delta_{\mathrm{S}}$. The accumulated phases can be dynamically calibrated to reach identical values using the pulses schematically plotted in Fig. 7(b). The time evolution is split into the following four-stage process. At stage 1, the array evolves naturally under the grid vector $\Lambda$. An $X$ pulse is applied to qubit 3 before stage 2. This transforms the grid vector into $X_{3}\Lambda X_{3}$ and reverses the velocity components of bond ‘W’ and ‘S’. Hence the accumulated phase for bond ‘W’ and ‘S’ seems a decrease while the phases increase for ‘N’ and ‘W’. This is then followed by an $X_{4}$ pulse, transforming the grid vector for the third stage into $X_{3}X_{4}\Lambda X_{3}X_{4}$ and reversing the velocities for bond ‘E’ and ‘W’. An $X_{4}$ pulse is then applied and transforms the grid vector for the final stage as $X_{4}\Lambda X_{4}$. The time-integrated phases for the four stages are shown in the bar charts Fig. 7(c.1-4). Assuming that the target phase is $\phi$ for all the bonds, the time durations $\tau_{0}$ to $\tau_{3}$ of the four stages can be solve by
@@
Applying our protocol to the most general case, we wish to find an appropriate pulse scheme to calibrate the bond strength for any quantum dot array with arbitrary connection. Apparently, if the array is loopless, such as the ones in Fig. 2, it is possible to reverse each bond independently. A pulse scheme to calibrate all bonds can be easily constructed. If the array topology involves loops, such as the rectangular grid example in Fig. 7, it is generally impossible to reverse all bonds independently. There, one must adopt combinations of different bond reversal schemes to find consistent solutions for the pulse locations and intervals. Here we prove that there is always an appropriate pulse scheme for the fully connected array. Note that this automatically indicates that an array with arbitrary connection can be calibrated as well.
To study the Pauli group conjugations $\{Q\Lambda Q\}$, it suffices to only consider assigning $X$ gates to the fully connected dot array. A particular assignment of $X$ gates, flips the bond velocities $(\Delta_{1},\Delta_{2},\Delta_{3},\cdots)$ according to an assignment vector, e.g, $\bm{a}_{k}=(+1,-1,+1,\cdots)^{\mkern-1.0mu\mathsf{T}}$ with $\lvert\bm{a}_{k}\rvert=n(n-1)/2\equiv N_{b}$ for a a fully connected array of $n$ dots. There are $2^{n}$ distinctive ways of assigning $X$ gates to the vertices, but different assignments can lead to same bond reversal scheme. We assume that there are at least $N_{a}$ linearly independent assignment vectors. The pulses intervals are solved in an equation similar to Eq. (75),
@@ IV.4 Tunable connectivity with SOI
The typical methods to switch on-and-off the interaction between neighboring spins require either biasing the dot chemical potentials or lowering the tunneling barrier between neighboring dots. In essence, these methods are can be viewed as the modulation of the exchange coupling energy $J$. This approach works well for two qubit systems, as verified by recent progresses reporting high fidelity figures. However, there are two problems that can hinder the wider application of such approach. First, the interdot coupling cannot be completely cut-off. Experimentally, $J$ is modulated over several order of magnitudes to distinguish the “on” and “off” state. The residue coupling at “off” state will result in errors that accumulate through the circuit. Second, such method does not scale well in terms of the control resources. Modulating the voltage of one particular gate electrode necessarily impact the the exchange coupling of all surrounding dots. As a result, the voltages on many gate electrodes must be simultaneous adjusted. The knowledge of how all the gates should synchronously change is obtained through prior calibration stage. One can imagine the tremendous efforts required to calibrate and accurately control the interdot coupling in a quantum chip with millions of qubits. A viable solution to this issue is to introduce tunable couplers between dots, such as superconducting Josephson junctions or transmon qubits [64]. Here we propose a whole new approach to control interdot coupling based on the theory developed in this work.
@@
One key indigent in our theory is the inclusion of spin-orbital coupling. This effect result in an additional spin-flipping channeling when tunneling among dots. However, as suggested by Eq. (36), these two channels act destructively to constitute an effective bond velocity
$\Delta_{w}=T_{w}-S_{w}=\frac{1}{2}J_{w}\left(\lvert t_{w}\rvert^{2}-\lvert s_{w}\rvert^{2}\right).$ (79)
@@
The typical methods to control inter-qubit coupling are centered around controlling the exchange energy $J_{w}$. Here we see that it is also possible to control the effective velocity by adjusting the ratio between $\lvert t_{w}\rvert$ and $\lvert s_{w}\rvert$. In particular, when $\lvert t_{w}\rvert=\lvert s_{w}\rvert$, no phase difference can accumulate across the bond and the bond can be seen as completely cut off. We note that the effective bond velocity can be interpret as an effective exchange energy that can take on negative values, in which the energy relation between spin singlet and triplet states are reversed. Such phenomenon have been reported in studies on few-electron coupled quantum dots [65, 66].
We can distinguish between the coupling and connectivity of two dots. Spin on two dots is coupled if the exchange energy is non-zero, while they are considered connected only if the effective bond velocity is non-zero. By tunning the relative strength between $\lvert t_{w}\rvert$ and $\lvert s_{w}\rvert$ while keeping the exchange energy finite, one can control the connectivity with always-on coupling.
@@
Figure 8: Dependence of the spin-conserved and spin-flipped tunneling fraction $\lvert t\rvert^{2}$ and $\lvert s\rvert^{2}$ on the spin-orbit length $x_{\mathrm{SO}}$ for a typical quantum dot array system. The bond connectivity can be controlled by tunning the relative strength between spin-conserved and spin flipped channel. In the subsets, we considered tuning the SOI strength for the right bond in a three-qubit array. We identity the (a) “on” status and (b) “off” status of this bond based on the effective bond velocity. This change of status can be achieved even though the coupling energy $J$ remains unchanged.
@@
Experimentally, both $t_{w}$ and $s_{w}$ are dependent on the SOI strength and magnetic field angle on the bond $w$. It is known that the SOI strength can be continuously tuned by varying the electric field strength. Hence for static and global the magnetic field, it is possible to control the SOI strength of a particular bond by changing the local electric field. Extra capacitor structure may be involved to suppress crosstalk and enable such precise control. Current spin qubit devices demonstrating such gate tunabelity is in [67] and also in [68]. For small SOI coupling $x_{\mathrm{SO}}\gg 2d$, the interdot tunneling is dominated by the spin-conserved process. As the SOI strength increases ($x_{\mathrm{SO}}$ decreases), $\lvert s\rvert^{2}$ picks up and will come across the decreasing $\lvert t\rvert^{2}$ at some point. This cross point always exists for magnetic field angle $\pi/4<\theta_{\mathrm{B}}<3\pi/4$ and define the “off” state of the relevant bond. In Fig. 8, we plotted the dependence of $\lvert t\rvert^{2}$ and $\lvert s\rvert^{2}$ on the spin-orbit length $x_{\mathrm{SO}}$, for a system with interdot spacing $1002d=$ 100 nm, characteristic Bohr radius $x_{0}=10$ nm and the angle between magnetic field and SOI vector $\theta_{\mathrm{B}}=\pi/3$. Assuming that we are varying the SOI strength for the second bond in a three-qubit array. The insets in Fig. 8 show the intrinsic gates for the (a) “on” state and (b) “off” state of the second bond, where we use parallel lines to indicate a open bond with finite velocity and cross lines to indicate closed bond with zero velocity. In state (a), the effective velocities for the two bonds are identical $\Delta_{1}=\Delta_{2}$, and the qubits experience a CZZ gate after time evolution $\tau=\pi/\Delta_{1}$. While in state (b), the same time evolution only induces a CZ gate between the left two qubits—as if the second bond is disconnected. This property holds for general many-qubit arrays. A particular bond can appear disconnected if the spin-conserved and spin-flipped tunneling strength through that bond are equal.

### v2

@@ IV.3 Dynamical calibration: general theory
We now present a general theory for the pulsed evolution in the presence of bond inhomogeneity. For an array of quantum dots with given connection, the intrinsic gate can be decomposed into the product of MQCP gates. The dynamics rule of the gate reduces to the time evolution in a high-dimensional phase spaces according the equation
@@
where $\bm{\Delta}=(\Delta_{1},\Delta_{2},\cdots)$ is the vector of the effective strength for all the connecting bonds for $\Delta_{w}=T_{w}-S_{w}$, and $\bm{\theta}=(\theta_{1},\theta_{2},\cdots)$ is the vector of target phase factors. The $\pi$-modulus implies that the target gate is represent by a high-dimensional lattice of target points. To implement simultaneous multi-qubit controlled-Z gates, all components of $\bm{\theta}$ are equal to $\pi$ and it is desirable that the effective strength $\Delta_{w}$ for all the participating bonds are all equal. In practice, the bond strength may differ from ideal values, resulting in unwanted errors in gate implementations.
We seek to dynamically calibrate the bond strengths into a uniform value. This is done by keeping the dots coupled while inserting single-qubit pulses during the evolution. In general, we can write the qubit-frame time evolution operator for the pulsed evolution as
@@
where the pulse $P_{n}$ is applied at time instance $\tau_{0}+\tau_{1}+\cdots+\tau_{n}$ to some vertex point, within the total duration $\tau=\tau_{0}+\tau_{1}+\cdots+\tau_{N}$. The reverse rotation $\mathrm{e}^{\mkern 1.0mu\mathrm{i}\mkern 1.0mu\tau H_{0}}$ in Eq. (73) transforms the lab-frame map to the qubit-frame. The $P_{n}$ gates are chosen from the Pauli group and they compose into $P_{N}P_{N-1}\cdots P_{1}=Q$. To proceed, we introduce a set of auxiliary gates $\{Q_{n}\}$ according to $Q_{n}=P_{n}P_{n-1}\cdots P_{1}$, with $Q_{0}=I$ and $Q_{N}=Q$. The time evolution can be rewritten as
@@
where we have invoked the approximation Eq. (63) for each time fragment in Eq. (74). Since $H_{0}$ involves only Pauli-Z operators $\{\sigma_{j}^{\mathrm{Z}}\}$, conjugation of $H_{0}$ by $Q_{n}$ only selectively flips the signs for the $\sigma_{j}^{\mathrm{Z}}$ operators and the resulting matrix is still diagonal in the computational basis. Hence we can commute and combine terms in Eq. (74) and further obtain
@@
The time evolution now comprises of three parts. First, an extra Pauli gate $Q$ is introduced by the applied pulses. It can be combined with latter gates in the circuit. Alternatively, we can simply choose a pulse sequence composing into $Q=I$. The second part in Eq. (75) is a set of local phase shifts induced by the $X_{j}$ and $Y_{j}$ conjugations with $\sigma_{j}^{\mathrm{Z}}$. Assuming $Q_{n}=\bigotimes_{j}Q_{nj}$ with $Q_{kj}\in\{I,X,Y,Z\}$ for the $j$th qubit in the array. For a given pulse scheme, the additional phases is explicitly determined by
@@
where $\operatorname{sig}(I)=\operatorname{sig}(Z)=1$ and $\operatorname{sig}(X)=\operatorname{sig}(Y)=-1$. The extra local phase shifts can be easily corrected with virtual-$Z$ gates and we may ignore them. The third part in Eq. (75) describes the entangling dynamics of the array, which is a simple product of $N+1$ time-independent stages. Each stage is generated by a Pauli-conjugated grid vector $\Lambda^{(n)}\equiv Q^{\dagger}_{n}\Lambda Q_{n}$ that can be decomposed as the Kronecker sum
@@
By expressing each bond vector as the diagonal matrix $\Lambda_{w}\mathrel{\widehat{=}}\operatorname{diag}(S_{w},T_{w},T_{w},S_{w})$, one can directly verify that $X$ or $Y$ conjugations on either dot exchange the $T_{w}$ and $S_{w}$ tunneling coefficients, while $\Lambda_{w}$ is invariant under $Z$ conjugations. Applying $X$ or $Y$ conjugation to both $j$ and $k$ exchanges the tunneling coefficients twice and leaves the bond vector in its original state. For example, $X_{j}\Lambda_{w}X_{j}=Z_{j}Y_{k}\Lambda_{w}Z_{j}Y_{k}\mathrel{\widehat{=}}\operatorname{diag}(T_{w},S_{w},S_{w},T_{w})$ and $X_{j}Y_{k}\Lambda_{w}X_{j}Y_{k}=\Lambda_{w}$. The exchange of tunneling coefficients effectively results in a reversal of velocity component for bond $w$, $\Delta_{w}\to-\Delta_{w}$. For the entire array of dots, an $X$ or $Y$ gate applied to one vertices point will simultaneously reverse the velocities for all the bonds connected to that dot. Combined with the geometry of the array, each $Q_{n}$ specifies the signs for all the bond velocity components. The $\Lambda^{(n)}$ at different stages of evolution are just different combinations of bond velocity components to be integrated with evolution time. This provide a method to dynamically calibrating the bond strength by properly choosing the time and locations of applied pulses.
Proper arrangement of the control pulses is the central to dynamic calibration. The arrangement of pulses is further determined by the connectivity of the array. If the array is loopless, i.e., the connecting bonds do not form a loop, such as the ones in Fig. 2, it is apparently possible to reverse each bond independently. A pulse scheme to calibrate all bonds can be easily constructed. If the array topology involves loops, it is impossible to reverse all bonds independently. There, one must adopt combinations of different bond reversal schemes to find consistent solutions for the pulse locations and intervals.
@@
To demonstrate dynamical calibration under sophisticated connectivity, we consider the rectangular grid example shown in Fig. 7(a). The dots are connected with four inhomogeneous bonds labeled by ‘N’, ‘E’, ‘W’ and ‘S’. Without loss of generality, we assume $\Delta_{\mathrm{N}}<\Delta_{\mathrm{E}}<\Delta_{\mathrm{W}}<\Delta_{\mathrm{S}}$. The accumulated phases can be dynamically calibrated to reach identical values using the pulses schematically plotted in Fig. 7(b). The time evolution is split into the following four-stage process. At stage 1, the array evolves naturally under the grid vector $\Lambda$. An $X$ pulse is applied to qubit 3 before stage 2. This transforms the grid vector into $X_{3}\Lambda X_{3}$ and reverses the velocity components of bond ‘W’ and ‘S’. Hence the accumulated phase for bond ‘W’ and ‘S’ seems a decrease while the phases increase for ‘N’ and ‘W’. This is then followed by an $X_{4}$ pulse, transforming the grid vector for the third stage into $X_{3}X_{4}\Lambda X_{3}X_{4}$ and reversing the velocities for bond ‘E’ and ‘W’. An $X_{4}$ pulse is then applied and transforms the grid vector for the final stage as $X_{4}\Lambda X_{4}$. The time-integrated phases for the four stages are shown in the bar charts Fig. 7(c.1-4). Assuming that the target phase is $\phi$ for all the bonds, the time durations $\tau_{0}$ to $\tau_{3}$ of the four stages can be solve by
@@
In order for the dynamical calibration protocol to work in the most general cases, we wish to find an appropriate pulse scheme to calibrate the bond strengths for any quantum dot array with arbitrary connectivity. Here we prove that there is always an appropriate pulse scheme for the fully connected array. Note that this automatically indicates that an array with arbitrary connection can be calibrated as well, as the later is only a subset of the former. It suffices to only consider assigning $X$ gates to the fully connected dot array. A particular assignment of $X$ gates, flips the bond velocities $(\Delta_{1},\Delta_{2},\Delta_{3},\cdots)$ according to an assignment vector, e.g, $\bm{a}_{k}=(+1,-1,+1,\cdots)^{\mkern-1.0mu\mathsf{T}}$ with $\lvert\bm{a}_{k}\rvert=n(n-1)/2\equiv N_{b}$ for a a fully connected array of $n$ dots. There are $2^{n}$ distinctive ways of assigning $X$ gates to the vertices, but different assignments can lead to same bond reversal scheme. We assume that there are at least $N_{a}$ linearly independent assignment vectors. The pulses intervals are solved in an equation similar to Eq. (78),
@@ IV.4 Tunable connectivity with SOI
The typical methods to switch on-and-off the interaction between neighboring spins require either biasing the dot chemical potentials or lowering the tunneling barrier between neighboring dots. In essence, these methods can be viewed as the modulation of the exchange coupling energy $J$. This approach works well for two qubit systems, as verified by recent progresses reporting high fidelity figures. However, there are two problems that can hinder the wider application of such approach. First, the interdot coupling cannot be completely cut-off. Experimentally, $J$ is modulated over several order of magnitudes to distinguish the “on” and “off” state. The residue coupling at “off” state will result in errors that accumulate through the circuit. Second, such method does not scale well in terms of the control resources. Modulating the voltage of one particular gate electrode necessarily impact the the exchange coupling of all surrounding dots. As a result, the voltages on many gate electrodes must be adjusted simultaneously. The knowledge of how all the gates should synchronously change is obtained through prior calibration stage. One can imagine the tremendous efforts required to accurately calibrate and control the interdot coupling in a quantum chip with millions of qubits. A viable solution to this issue is to introduce tunable couplers between dots, such as superconducting Josephson junctions or transmon qubits [69]. Here we propose a whole new approach to control interdot coupling based on the theory developed in this work.
@@
One key indigent in our theory is the inclusion of SOI. This effect result in an additional spin-flipping channeling when tunneling among dots. However, as suggested by Eq. (38), these two channels act destructively to constitute an effective bond strength
$\Delta_{w}=T_{w}-S_{w}=\frac{1}{2}J_{w}\left(\lvert\tilde{t}_{w}\rvert^{2}-\lvert\tilde{s}_{w}\rvert^{2}\right).$ (82)
@@
The typical methods to control inter-qubit coupling are centered around the modulation of $J_{w}$. Here we see that it is also possible to control the effective strength by adjusting the ratio between $\lvert\tilde{t}_{w}\rvert$ and $\lvert\tilde{s}_{w}\rvert$. In particular, when $\lvert\tilde{t}_{w}\rvert=\lvert\tilde{s}_{w}\rvert$, no phase difference can accumulate across the bond and the bond can be seen as completely cut off. We note that $\Delta_{w}$ can be interpret as an effective exchange energy that can take on negative values, in which the energy relation between spin singlet and triplet states are reversed. Such phenomenon have been reported in studies on few-electron coupled quantum dots [70, 71]. We can hence distinguish between the “coupling” and “connectivity” of two dots. Spin on two dots is coupled if the exchange energy $J_{w}$ is non-zero, while they are considered connected only if the effective coupling strength $\Delta_{w}$ is non-zero. By tunning the relative strength between $\lvert\tilde{t}_{w}\rvert$ and $\lvert\tilde{s}_{w}\rvert$ while keeping the exchange energy finite, one can control the connectivity with always-on coupling.
@@
Figure 8: Dependence of the spin-conserved and spin-flipped tunneling fraction $\lvert\tilde{t}\rvert^{2}$ and $\lvert\tilde{s}\rvert^{2}$ on the spin-orbit length $x_{\mathrm{SO}}$ for a typical quantum dot array system. The bond connectivity can be controlled by tunning the relative strength between spin-conserved and spin flipped channel. In the subsets, we considered tuning the SOI strength for the right bond in a three-qubit array. We identity the (a) “on” status and (b) “off” status of this bond based on the effective bond strength. This change of status can be achieved even though the coupling energy $J$ remains unchanged.
@@
Experimentally, both $\tilde{t}_{w}$ and $\tilde{s}_{w}$ are dependent on the SOI strength and magnetic field angle on the bond $w$. It is known that the SOI strength can be continuously tuned by varying the electric field strength. Hence for static and global the magnetic field, it is possible to control the SOI strength of a particular bond by changing the local electric field. Extra capacitor structure may be involved to suppress crosstalk and enable such precise control. Current spin qubit devices demonstrating such gate tunability is in [72] and also in [73]. For small SOI coupling $x_{\mathrm{SO}}\gg 2d$, the interdot tunneling is dominated by the spin-conserved process. As the SOI strength increases ($x_{\mathrm{SO}}$ decreases), $\lvert\tilde{s}\rvert^{2}$ picks up and will come across the decreasing $\lvert\tilde{t}\rvert^{2}$ at some point. This cross point always exists for magnetic field angle $\pi/4<\vartheta_{\mathrm{B}}<3\pi/4$ and define the “off” state of the relevant bond. In Fig. 8, we plotted the dependence of $\lvert\tilde{t}\rvert^{2}$ and $\lvert\tilde{s}\rvert^{2}$ on the spin-orbit length $x_{\mathrm{SO}}$, for a system with interdot spacing $1002d=$ 100 nm, characteristic Bohr radius $x_{0}=10$ nm and the angle between magnetic field and SOI vector $\vartheta_{\mathrm{B}}=\pi/3$. Assuming that we are varying the SOI strength for the second bond in a three-qubit array. The insets in Fig. 8 show the intrinsic gates for the (a) “on” state and (b) “off” state of the second bond, where we use parallel lines to indicate a connected bond and cross lines to indicate disconnect bond. In state (a), the effective coupling strengths for the two bonds are identical $\Delta_{1}=\Delta_{2}$, and the qubits experience a CZZ gate after time evolution $\tau=\pi/\Delta_{1}$. While in state (b), the same time evolution only induces a CZ gate between the left two qubits while leaving the third qubit intact. This property holds for general many-qubit arrays. A particular bond can appear disconnected if the spin-conserved and spin-flipped tunneling strengths are equal.